OpenStudy (anonymous):

What is the next term in this sentence? 2,6,10,14

OpenStudy (jhannybean):

What is the common difference?

sammixboo (sammixboo):

Well, look at the first two. 2, 6 What did they do to 2 to get 6?

OpenStudy (jhannybean):

\[d=a_2-a_1\]

OpenStudy (anonymous):

it is 18 right??

sammixboo (sammixboo):

Right :)

OpenStudy (jhannybean):

\[\checkmark\]

OpenStudy (anonymous):

thx :P

sammixboo (sammixboo):

2 + 4 = 6 6 + 4 = 10 10 + 4 = 14 14 + 4 = 18

OpenStudy (anonymous):

thx so much lol